# Variational autoencoders stabilise TCN performance when classifying weakly labelled bioacoustics data

### v1

@@ Abstract
Passive acoustic monitoring (PAM) data is often weakly labelled, audited at the scale of detection presence or absence on timescales of minutes to hours. Moreover, this data exhibits great variability from one deployment to the next, due to differences in ambient noise and the signals across sources and geographies. This study proposes a two-step solution to leverage weakly annotated data for training Deep Learning (DL) detection models. Our case study involves binary classification of the presence/absence of sperm whale (Physeter macrocephalus) click trains in 4-minute-long recordings from a dataset comprising diverse sources and deployment conditions to maximise generalisability. We tested methods for extracting acoustic features from lengthy audio segments and integrated Temporal Convolutional Networks (TCNs) trained on the extracted features for sequence classification. For feature extraction, we introduced a new approach using Variational AutoEncoders (VAEs) to extract information from both waveforms and spectrograms, which eliminates the necessity for manual threshold setting or time-consuming strong labelling. For classification, TCNs were trained separately on sequences of either VAE embeddings or handpicked acoustic features extracted from the waveform and spectrogram representations using classical methods, to compare the efficacy of the two approaches. The TCN demonstrated robust classification capabilities on a validation set, achieving accuracies exceeding 85% when applied to 4-minute acoustic recordings. Notably, TCNs trained on handpicked acoustic features exhibited greater variability in performance across recordings from diverse deployment conditions, whereas those trained on VAEs showed a more consistent performance, highlighting the robust transferability of VAEs for feature extraction across different deployment conditions.
keywords:
Weak labelling, Bioacoustics, Temporal Convolutional Networks, Variational Autoencoders
1 Introduction
This study addresses the challenge of working with the complexities of weakly labelled bioacoustics datasets, with a specific focus on sperm whale clicks as a case study. Sperm whales (Physeter macrocephalus) emit frequent and easily detectable clicks for the majority of their dive time. Maximum source levels of 240dB re 1 µPa have been recorded for these clicks [1], with click lengths of 1ms and average inter-click intervals of around 0.5 seconds between regular vocalisations [2]. There regularity of these clicks makes Passive Acoustic Monitoring (PAM) a particularly suitable technique to study sperm whales. PAM is a method used to survey and monitor wildlife and environments through the deployment of acoustic recorders. In the case of marine animals, PAM offers the advantage of enabling continuous data collection, even under adverse weather conditions or when visual surveys are unfeasible. The analysis of acoustic recordings is frequently used to study sperm whale distribution in space and time, and because sperm whales are widely distributed there are dozens of underwater datasets featuring their vocalisations. The acoustic and spectral properties of individual sperm whale clicks contain limited information to distinguish them from other transients, but their regularity in terms of amplitude and inter-click-intervals provides crucial contextual information. Therefore, it is desirable for a detection system to incorporate both local and contextual information in order to identify click trains.
The primary objective of this work was to develop a robust classifier capable of distinguishing between multiple minute-long audio segments containing sperm whale clicks and those without. To ensure diversity of training data, we used multiple datasets, allowing the model to focus on learning vocalisation characteristics robust to changes in recording equipment and/or environment. Additionally, we tested various feature extraction approaches to compensate for the relative scarcity of well-labelled data for individual vocalisations.
@@ 1 Introduction
As our capacity for collecting large quantities of acoustic data continues to grow, so does the demand for automated detection systems to effectively process and analyse it. While PAM has traditionally relied upon many different computational techniques for data analysis, such as spectrogram correlation [3], random forests [4], and support vector machines [5], Deep Learning (DL) methods have gained popularity for tasks like the detection of animal sounds from large acoustic datasets [6].
Underwater acoustic datasets pose a unique set of challenges when it comes to training deep learning methods. Some of these challenges arise from the substantial cost associated with data collection and auditing, which is often aggravated by the scarcity of events in the datasets. As a result, the sizes of annotated datasets and the number of examples of each signal of interest utilised in these studies are considerably smaller, often orders of magnitude below the dimensions of typical benchmark datasets. For instance, the DCLDE 2024 acoustic dataset, a specialized dataset used for developing and benchmarking algorithms for detecting, classifying, localizing, and estimating the density of marine mammal calls, contains only a few hundred labeled examples of each call type [7]. Similarly, the Xeno-Canto dataset [8], likely the largest and most used focal bird song database, contains around 678,000 weakly labelled audio recordings of 10,300 species, significantly smaller in scale compared to other benchmark datasets like ImageNet, which contains over 14 million labeled images [9].
Another challenge is the fact that most underwater PAM data is annotated as presence or absence of vocalisations within time periods on scales of minutes to hours, rather than at a resolution of single vocalisations, a practice known as weak labelling. This is often because PAM data is collected on duty cycles to maximise deployment times without over-sampling. Common approaches for annotating bioacoustic datasets include auditing several minutes of data sampled at regular intervals and marking the presence of different species vocalisations or specifying the start and end times of acoustic events, when such events may include intermittent periods of silence [10]. This method is favoured for its cost-effectiveness and real-time feasibility during field surveys. Moreover, it reduces the degree of interpretation required and, consequently, the variability among labels involved in the annotation process. However, this approach significantly increases computational demands by generating data that is larger and more complex than that produced by strong label approaches, which in turn requires larger and more complex models to process it, often rendering it practically infeasible.
This study addresses two of the most prevalent challenges in training deep DL models with bioacoustics data. Firstly, we tackle the issue of small dataset sizes and the high variability in ambient noise and acoustic data across different sources and geographical regions by combining data from multiple sources collected under various deployment conditions. Secondly, we address the problem of weak labeling by proposing a workflow that leverages existing weakly annotated data to train robust DL classification models.
In this pursuit, we propose a novel approach based on unsupervised variational autoencoders (VAEs) [11] to learn a low-dimensional feature representation space, while at the same time eliminating the need for a manual threshold setting or strong labels for individual target sound events. We assessed the ability of VAEs to learn meaningful patterns and extract information from waveforms and spectrogram representations without prior knowledge of the target sounds, and we compare them to classical methods used in earlier works to extract acoustic parameters, including root-mean-square calculations, energy level sums, and spectral properties [12].
After data feature extraction, we used Temporal Convolutional Networks (TCNs), a deep learning architecture tailored for sequence processing [13], to classify acoustic recordings based on the extracted feature patterns. While TCNs have been successfully employed for audio data classification in recent years [13, 14, 15], the present study is a novel application of these methods in the bioacoustics field used in conjunction with unsupervised data representation via VAEs.
1.1 Main innovations and contributions
The aims and contributions of this study were as follows:
(1) We collated PAM recordings from various sources and deployment conditions into a comprehensive dataset that includes different types of background noise and other vocalising species to promote generalisation of the solution. This dataset is weakly labelled with annotations for periods of multiple minutes, and containing well-balanced instances of the signal of interest (i.e., sperm whale vocalisations) and various noise types.
(2) We trialled a new approach incorporating VAEs to extract acoustic features from concise audio segments without requiring additional, time-consuming and thus costly, labelling. Furthermore, we assessed the data representation efficacy of the extracted features against conventionally employed handcrafted acoustic features.
(3) We incorporated a 4-min long temporal context into the classification process by using TCNs trained on extracted features for sequence classification.
(4) We investigated the effect of annotation length on the different models’ performance to determine whether a greater emphasis on temporal context, or higher temporal resolution in extracted features, helps obtain better results in the detection task.
2 Related work
This section provides an overview of the key challenges and exiting solutions in bioacoustics data processing for training DL models, as well as on existing sperm whale detection methods. The first section offers a review of the existing publicly accessible marine bioacoustics datasets. The second section addresses the particularities of bioacoustics data, focusing on dataset sizes and weak labelling. The third section reviews commonly used feature engineering methods for acoustic data. The fourth section reviews section reviews commonly used deep learning sequential classification methods for bioacoustics data. Finally, the fifth and last section provides an overview of the state-of-the-art in acoustic detectors for sperm whale clicks.
2.1 Marine bioacoustics datasets
In marine bioacoustics, most of the public datasets available have been produced as part of the DCLDE workshop series [16], with the last four workshops having data publicly available at the time of writing this manuscript [17, 18, 19, 7].
While the DCLDE workshops as collaborative initiatives are highly valuable and establish a solid baseline, each dataset is derived from one or two related surveys, utilizes only a single type of recording equipment, and covers a small geographical and temporal range. To build highly general models that can perform well across diverse conditions, baseline training sets must comprise data from multiple domains and varied deployment conditions [20]. Thus, there is a significant gap in the availability of such comprehensive datasets, which this work aims to address.
2.2 Handling weakly labelled datasets
Many open-source annotated audio datasets predominantly provide weak labels [21], a case that is specially true for bioacoustics datasets [8, 19]. The challenge of dealing with weak labels in audio classification has garnered significant attention across various domains [21, 22, 23]. A straightforward approach involves discarding temporal context and representing audio signals as a long-term statistical distribution of their short-term spectral features, commonly known as the bag of frames approach, extensively used in music [24, 25] and soundscape [26] recognition. However, the efficiency of this method in analysing complex, dynamic sounds is limited [27], and hence context-aware approaches have been needed to achieve performance improvements in these domains [28]. Similarly, the training approach known as the ’strong label assumption’ segments the audio into frames and assigns a single parent label to each frame [29]. This results in an increase in the number of smaller training instances, which can be advantageous for smaller datasets. However, there is a risk that some of these instances may be incorrectly labelled.
Other strategies for handling weakly labelled audio datasets vary in the degree of expert involvement required. These range from further expert annotation [30] or active learning methods [31] to semi-automatic annotation techniques, like combining automatic clustering with manual annotation [32, 33, 34, 35], and the adoption of fully independent labeling functions such as joint-detection-classification models, attention methods, and label pooling [36, 37, 38]. However, while all these methods strive to identify the specific events that correspond to the annotated class, they do not fully incorporate all the temporal context available within the long interval into the decision-making process.
2.3 Feature engineering
Classic methods for the automatic detection of bioacoustic sounds often rely on the generation of specific acoustic parameters from waveform and spectrogram representations [39, 40]. These acoustic parameters can range from simple energy sums to complex engineered features like the energy fingerprint [41]. Some of the parameters most commonly used to detect and classify odontocete clicks include peak and mean frequencies, spectral shape and distribution, signal-to-noise ratio, duration, inter-click-intervals [42, 43, 12, 44, 35]. Recently, Li et al. [41] obtained high accuracies in species classification using a complex feature that combines energy and temporal information. Notably, Frasier [45] proposed using acoustic indices to cluster data into classes that are then used to create a labeled dataset to train DL models. However, these indices are often very dependant on site and situation, and can lead to poor performance on new previously unheard soundscapes [46].
An alternative method for feature engineering involves utilising unsupervised learning approaches such as autoencoders [47]. Autoencoders have been demonstrated to produce representations surpassing handcrafted acoustic features across various analysis tasks in animal calls, including representing acoustic similarity and multi-species discrimination [48, 49]. Reeves Ozanich et al. [50] successfully applied autoencoders in conjunction with a clustering method to differentiate between whale and fish calls. However, it remains unclear whether autoencoder representations actually outperform acoustic features in extracting meaningful features for complex downstream tasks as a direct comparison has not been done.
Furthermore, autoencoders have some limitations, including a lack of a probabilistic framework that may not produce a continuous latent space, limiting their ability to capture complex data distributions and generalise to new data points. In contrast, Variational AutoEncoders (VAEs), encode input data into parameters of a normal distribution within the latent space, enhancing robustness to variations and noise in the input data [51]. This approach also facilitates smooth interpolation between data points in the continuous latent space, allowing for meaningful exploration and manipulation of data representations.
2.4 Sequential classification
Various neural network architectures based on sequential data processing have been developed and applied on bioacoustics datasets. Recurrent Neural Networks [52] and Long Short-Term Memory networks [53] are commonly used for tasks like sound event detection and species identification [54, 55, 56]. Some efforts have also focused on combining methods for high-resolution feature extraction with sequencing models. For instance, Madhusudhana et al. [55] utilised a pairing of a Convolutional Neural Network (CNNs) and Recurrent Neural Network to detect fin whale calls, achieving high performance metrics. This approach was also used by Wang et al. [57] and Winsen et al. [58] for detecting calls of land mammals. In the field of music informatics, Won et al. [59] combined a CNN with self-attention layers for music genre classification. Notably, all these studies have leveraged strong labelled datasets for training the feature extraction CNN.
Recently, Temporal Convolutional Networks (TCNs) have emerged as a novel approach to incorporate temporal context hierarchically into pattern recognition, showing promising results in sequence modelling tasks [13]. Recent research has explored the application of TCNs in bioacoustics, particularly for species that vocalise in characteristic bouts [6].
2.5 Sperm whale detectors
A suite of algorithms have been developed to detect sperm whale echolocation clicks. Some approaches used include time-frequency analysis of the clicks [42, 60], energy comparisons [43], or the use of the Teager-Kaiser Energy Operator [61]. Recently, studies have used CNNs to distinguish between spectrograms containing or not containing odontocete [62] or more specifically sperm whale [63] clicks. However, all of these studies focus on detections of individual clicks, rather than incorporating multiple clicks into the decision process.
@@ 2.5 Sperm whale detectors
Recent studies have attempted to address the task of detecting sperm whale click trains through heuristic methods, including the work of Macaulay [64] for detecting porpoise click trains, later applied to sperm whales in Webber et al. [65]. Another approach used deep learning models to classify long sequences (4 minutes) based on the sequence of click detections on shorter (5 second) windows [66]. While both approaches were successful in their respective tasks, the method proposed by Macaulay [64] requires extensive manual fine-tuning to suit specific datasets, and the deep learning approach of Garrobé et al. [66] is limited by the requirement for finely labeled data (in this case, annotations at a 5-second level), which is not always available.
3 Methods
Our proposed detection and classification method consists of two parts. In the first step the high resolution input data is mapped into a latent space without the need of any labels, and in the second step the resultant sequence of low-dimensional feature representations is used to predict the output class from the extracted features. This strategy leverages existing weak annotations to extract both acoustic features at a local scale and temporal patterns to classify audio segments. The development of the acoustic classifier involves a systematic workflow that can be divided into four different stages (Figure 1) described one by one next in the following subsections.
@@ 3 Methods
Figure 1: Schematic diagram outlining the methodology employed to develop an acoustic classifier from multiple sources of weakly labelled data. The four main steps followed for this study are (1) standardising annotations and recordings from different sources and creating two datasets that were representative of the variabilities in signal and noise, (2) developing and training different feature extraction methods, (3) training multiple recording classifiers for detecting sperm whale click trains based on temporal sequences of extracted features, and finally (4), comparing the performances of the trained classifiers.
3.1 Audio data annotation
@@
To enhance the robustness and generalisability of the solution, a comprehensive training set was constructed by combining data from various sources and recording environments. Specifically, this dataset was compiled from data collected by multiple studies utilising various methods and protocols, covering different geographical and temporal scales (Table 1). By collating data from various sources, we aim to introduce variabilities in the training and validation sets across various factors that would otherwise impact the classifier’s ability to transfer to unseen environments. These factors included:
@@
    The presence of anthropogenic and natural ambient sounds, as well as other vocalizing species, and signal-to-noise ratio of the recordings.
3.1.1 The 4-minute dataset
The main dataset utilized in this study comprises data from six distinct research studies conducted across various global locations, with data collection efforts in Atlantic, Southern, Arctic and Indian oceans (AS dataset, Webber et al. [65]), the Balearic Sea (BAL datasets, Garrobé et al. [66]), the California Blight (CAL dataset, DCLDE [17]), the Caribbean Sea (CS dataset, Vachon et al. [67]), the Icelandic off-shore waters (ICE dataset, Wensveen et al. [68]) and the Mediterranean Sea (MED dataset, Lewis et al. [69]). The BAL dataset was divided into its 4 deployments (BAL_1, BAL_2, BAL_3, BAL_4) given that there was enough data available and previous studies suggest high variability between deployments. The AS, CS and MED datasets were collected using towed hydrophones, while the BAL, CAL, CS and ICE data were collected using moored autonomous recorders. The detail and quality of the annotations also varied greatly from dataset to dataset, going from individual-level click labelling (MED dataset), to labelling at a 4-minute level (BAL and CS datasets), or semi-automatic labelling using automated acoustic algorithms to do an initial processing of the data (AS and ICE datasets).
@@ 3.1.1 The 4-minute dataset
Given the diverse recording conditions and labelling formats across data from different deployments, a necessary first step was standardising both recordings and labels. A uniform labelling format was established, which entailed dividing recordings into 4-minute segments that were then labelled either as positive (containing sperm whale clicks) or negative (not containing sperm whale clicks), derived from the original annotations. The 4-minute length was chosen due to it being the finest common resolution available across all sources. Additional information, such as the presence of other sound sources like ships and other vocalising animals, was retained when available. These annotations were solely kept as references to ensure a diverse range of noise sources in the final dataset but were not utilized as labels for training.
@@
After selecting the 4 minute segments and standardising the labels to either positive or negative for the whole segment, a pre-processing stage was implemented to standardise the recordings across diverse deployments. In cases of recordings featuring multiple channels, only the first channel was retained. Furthermore, any DC offset present in the signal was removed. Subsequently, all files were decimated to a uniform sampling rate of 48 kHz. To eliminate low frequency noise a 1-20kHz bandpass 6-pole Butterworth filter was applied, leaving a frequency range that captures the majority of the sperm whale click energy to analyse [71].
3.1.2 The 30-second dataset
@@
We constructed a dataset comprising 30-second audio files that served as the basis for examining the trade-off between increased temporal context for decision-making and shorter audio segments allowing for higher resolution features to be generated and extracted. This dataset was comprised of data from the AS and MED datasets, the two datasets where annotations were available at a single click level (Table 1).
3.2 Feature engineering
3.2.1 Extraction of clips for training feature engineering methods

### v2

# Temporal Feature Learning in Weakly Labelled Bioacoustic Cetacean Datasets via a Variational Autoencoder and Temporal Convolutional Network: An Interdisciplinary Approach

@@ Abstract
Bioacoustics data from Passive acoustic monitoring (PAM) poses a unique set of challenges for classification, particularly the limited availability of complete and reliable labels in datasets due to annotation uncertainty, biological complexity due the heterogeneity in duration of cetacean vocalizations, and masking of target sounds due to environmental and anthropogenic noise. This means that data is often weakly labelled, with annotations indicating presence/absence of species over several minutes. In order to effectively capture the complex temporal patterns and key features of lengthy continuous audio segments, we propose an interdisciplinary framework comprising dataset standardisation, feature extraction via Variational Autoencoders (VAE) and classification via Temporal Convolutional Networks (TCN). This approach eliminates the necessity for manual threshold setting or time-consuming strong labelling. To demonstrate the effectiveness of our approach, we use sperm whale (Physeter macrocephalus) click trains in 4-minute recordings as a case study, from a dataset comprising diverse sources and deployment conditions to maximise generalisability. The value of feature extraction via the VAE is demonstrated by comparing classification performance against the traditional and explainable approach of expert handpicking of features. The TCN demonstrated robust classification capabilities achieving AUC scores exceeding 0.9.
I Introduction
@@
Passive Acoustic Monitoring (PAM) is a method used to survey and monitor wildlife and environments using acoustic recorders. In the case of marine animals, PAM offers the advantage of enabling continuous data collection, even under adverse weather conditions or when visual surveys are unfeasible. As our capacity for collecting large quantities of acoustic data continues to grow, so does the demand for automated detection systems to effectively process and analyse it. While PAM has traditionally relied upon many different computational techniques for data analysis [32, 18, 11], Deep Learning (DL) methods have gained popularity for tasks that require detection of animal sounds from large acoustic datasets [44].
However, underwater PAM datasets pose a unique set of challenges when it comes to training DL methods, arising from the substantial cost associated with data collection and auditing, often aggravated by the scarcity of target events in the datasets. This results in considerably small annotated datasets with few instances of each signal of interest, both orders of magnitude below the dimensions of typical benchmark datasets [48, 53, 24]. Moreover, bioacoustics baseline datasets are often derived from limited surveys using single equipment types and narrow spatiotemporal ranges [8, 9, 10]. While these datasets are very valuable to establish a solid baseline for DL biacoustic studies, their reduced scope limits the generalisability of the models trained [36]. Finally, most underwater PAM data are annotated as presence/absence of vocalizations over periods of several minutes rather than individual calls, a practice known as weak labelling. Two common annotation approaches are auditing sampled intervals to mark species presence and marking acoustic event boundaries, which may include intermittent silence. These practices balance cost-effectiveness with field practicality, while also reducing variability between annotators [37]. However, weak labelling approaches significantly increase computational demands by generating data that is larger and more complex than that produced by strong labels. In turn, weak labelled datasets require larger and more complex models to process compared to finely annotated data, often rendering it practically infeasible.
One case for which weak labeling is especially fitting is sperm whale (Physeter macrocephalus) vocalizations. Sperm whales emit frequent and easily detectable clicks for the majority of their dive time. Maximum source levels of 240dB re 1 µPa have been recorded for these clicks [35], with click lengths of 1ms and average inter-click intervals of around 0.5 seconds between regular vocalisations [51]. The regularity of these clicks makes PAM a particularly suitable technique to study sperm whales, and because they are widely distributed, there are dozens of underwater datasets featuring their vocalizations. The acoustic and spectral properties of individual sperm whale clicks contain limited information to distinguish them from other transients, but their regularity in terms of amplitude and inter-click-intervals provides crucial contextual information. This makes weak labeling a practical choice for annotating sperm whale acoustic data, as it enables the annotator to leverage local information of individual clicks and broader temporal patterns into the decision. An effective detection system should therefore integrate both levels of spatiotemporal information to reliably identify click trains.
@@
A suite of algorithms have been developed to detect sperm whale echolocation clicks, including approaches based on the time-frequency analysis of the clicks [33, 34], energy comparisons [22], or the use of the Teager-Kaiser Energy Operator [19]. In the field of DL, there has been work using convolutional neural networks to distinguish between spectrograms containing or not containing odontocete [30] or, more specifically, sperm whale [3] clicks. However, all such studies focus on detections of individual clicks, rather than incorporating multiple clicks into the decision process, limiting capture of the full vocalizations. Recent studies have attempted to address the task of detecting sperm whale click trains through heuristic methods, including the work of Macaulay [31] for detecting porpoise click trains, later applied to sperm whales in Webber et al. [49]. Another approach used deep learning models to classify long sequences (4 minutes) based on the sequence of click detections on shorter (5 second) windows [14]. While both approaches were successful in their respective tasks, the method proposed by Macaulay [31] requires extensive manual fine-tuning to suit specific datasets, and the DL approach of Garrobé Fonollosa et al. [14] is limited by the requirement for finely labeled data (in this case, annotations at a 5-second level), which is not always available.
This study proposes a new interdisciplinary methodology for detecting sperm whale click trains that addresses two of the most prevalent challenges of working with the complexities of bioacoustics data and training DL models for acoustic detection. Firstly, we tackle the issue of small dataset sizes and high variability in environmental and anthropological noise and acoustic data across different sources and geographical regions by curating a dataset from multiple sources collected under various deployment conditions. Secondly, we address the problem of weak labeling by proposing a workflow that leverages existing weakly annotated data to train robust DL classification models. This comprises feature extraction via an unsupervised variational autoencoder (VAE) [20] that captures complex and temporal patterns, both locally and long-term with its generative nature allowing it to learn robust features from noisy and complex time-series data. This avoids the need for setting manual thresholds or providing strong labels. The extracted features are then classified via a Temporal Convolutional Network (TCN), a DL architecture tailored for sequence processing [1], that have been successfully employed for audio data classification in recent years [1, 7, 54] to capture long-term dependencies and temporal context. To the best of our knowledge, this paper is the first to propose the VAE-TCN framework in the context of bioacoustic data detection and classification.
Traditional bioacoustic detection often relies on hand-made parameters such as peak frequencies, spectral shape, and inter-click intervals [46, 45, 33]. While recent methods using energy-temporal features [27] or acoustic indices [12] show promise, they often fail in new soundscapes [42]. Autoencoders [5] and VAEs have been shown to surpass handcrafted features for tasks such as species discrimination [16, 40] and call clustering [38], but direct comparisons for complex tasks remain lacking. We assessed the ability of VAEs to learn meaningful patterns and extract information from waveforms and spectrogram representations without prior knowledge of the target sounds. In order to evaluate the learning efficacy of VAEs, we compare them to traditional, expert-led methods used in earlier works [6] to select acoustic parameters, including root-mean-square calculations, energy level sums, and spectral properties.
The contributions of this study can be summarized as follows:
  1. 1.
    Curation of PAM recordings from various sources and deployment conditions into a comprehensive dataset that includes different types of background noise and other vocalising species to promote generalisation of the solution. Recordings from different studies were re-sampled and cleaned, with labels standardised and erroneous annotations removed. This dataset is weakly labelled with annotations for periods of multiple minutes, and containing well-balanced instances of the signal of interest (i.e., sperm whale vocalizations) and anthropogenic and environmental noise.
  2. 2.
    Feature extraction via VAEs to learn acoustic features from audio segments or click trains without requiring additional, time-consuming and thus costly, labelling.
  3. 3.
    Classification on 4-min long vocalization sequences, via TCNs trained on the above VAE-extracted features.
  4. 4.
    Evaluation of the efficacy of the extracted features in terms of classification accuracy against conventionally employed handcrafted acoustic features.
  5. 5.
    Evaluating the importance of annotation length on the different models’ performance to determine whether a greater emphasis on temporal context, or higher temporal resolution in extracted features, influences performance.
II Methods
Our proposed framework follows a systematic workflow comprising four stages, as shown in FIG.  1. Each of the four stages is described next.
@@ II Methods
Figure 1: Schematic diagram outlining the methodology employed to develop an acoustic classification framework from diverse datasets of weakly labelled data. The four main steps followed for this study are (1) standardising annotations and recordings from different sources and creating two datasets that were representative of the variabilities in anthropogenic and environmental noise, (2) VAE-based feature extraction, (3) detection and classification of sperm whale click trains based on TCN, and finally (4), evaluation of the value of feature extraction vs handpicked, expert-led features and annotation length of the temporal sequence.
II.1 PAM dataset curation
@@
To enhance the robustness and generalisability of the solution, a comprehensive dataset was compiled from data collected by multiple studies utilising various sensors and recording protocols in different types of environment, covering different geographical and temporal scales (Table 1). By collating data from various sources, we aim to ensure robustness to the variabilities in the PAM datasets across various factors that would otherwise impact the classifier’s generalisation ability. These factors included:
@@
    The presence of anthropogenic and natural ambient sounds, as well as other vocalising species, and signal-to-noise ratio of the recordings.
  4. 4.
    Annotation variability from individual-level click labelling to 4-minute click train labelling to semi-automatic labelling.
@@
The main dataset used in this study comprises data from six distinct research studies conducted across various global locations, with data collection efforts in Atlantic, Southern, Arctic and Indian oceans (AS dataset, Webber et al. [49]), the Balearic Sea (BAL datasets, Garrobé Fonollosa et al. [14]), the California Blight (CAL dataset, DCLDE [8]), the Caribbean Sea (CS dataset, Vachon et al. [47]), the Icelandic off-shore waters (ICE dataset, Wensveen et al. [50]) and the Mediterranean Sea (MED dataset, Lewis et al. [26]). The BAL dataset was divided into its 3 deployments (BAL_1, BAL_2, BAL_3) given that there was enough data available and previous studies [15] suggest high variability between deployments. The AS, CS and MED datasets were collected using towed hydrophones, while the BAL, CAL, CS and ICE data were collected using moored autonomous recorders. The detail and quality of the annotations also varied greatly from dataset to dataset, going from individual-level click labelling (MED dataset), to labelling at a 4-minute level (BAL and CS datasets), or semi-automatic labelling using automated acoustic algorithms to do an initial processing of the data (AS and ICE datasets).
Given the diverse recording conditions and labelling formats across data from different deployments, a necessary first step was standardising both recordings and labels. A uniform labelling format was established, which entailed dividing recordings into 4-minute segments that were then labelled either as positive (containing sperm whale clicks) or negative (not containing sperm whale clicks), derived from the original annotations. The 4-minute length was chosen due to it being the finest common resolution available across all sources. Additional anthropological and environmental information, such as the presence of other sound sources like ships and other vocalising animals, was retained when available. These annotations were solely kept as references to ensure a diverse range of anthropological and environmental noise sources in the final dataset but were not used as labels for training.
@@
In addition to the above 4-minute segment dataset, a dataset comprising 30-second audio files was constructed that served as the basis for examining the trade-off between increased temporal context for decision-making and shorter audio segments allowing for higher resolution features to be generated and extracted. This dataset was comprised of data from the AS and MED datasets, the two datasets where annotations were available at a single click level (Table 1). The same pre-processing steps as in the 4-minute dataset were applied to the 30-second segments.
II.2 Pre-processing
After selecting the 4-minute and 30-second segments and standardising the labels to either positive or negative for the whole segment, a pre-processing stage was implemented to standardise the recordings across diverse deployments. In cases of recordings featuring multiple channels, only the first channel was retained. Furthermore, any DC offset present in the signal was removed. Subsequently, all files were decimated to a uniform sampling rate of 48 kHz. To eliminate low-frequency noise, the signal was filtered via a 1-20kHz bandpass 6-pole Butterworth filter, leaving a frequency range that captures the majority of the sperm whale click energy to analyse [17].